OpenStudy (anonymous):

How do you determine when to use addition or subtraction when solving an absolute value inequality using a graph?

OpenStudy (anonymous):

Example:

OpenStudy (anonymous):

I thought distance was always supposed to be negative, but some of the answers to my questions use a positive.

OpenStudy (lynfran):

no distance is always positive

OpenStudy (anonymous):

The answers use negatives.

OpenStudy (lynfran):

because we are talking about absolute values .... we usually get 2 answers one is positive and one is negative..

OpenStudy (lynfran):

@DecentNabeel wat do u think..?

OpenStudy (triciaal):

distance is always a positive unit of measure between 2 points

OpenStudy (lynfran):

for absolute .... example |x+3|=5 then (x+3)=+5 and (x+3)=-5

OpenStudy (triciaal):

one approach to doing absolute value problems is to split in 2 the positive and the negative solve each then find the combined solution

OpenStudy (anonymous):

I'm not solving an actual absolute value inequality, I am trying to write absolute value inequalities using a graph.

OpenStudy (triciaal):

read again slowly

OpenStudy (triciaal):

solve each and put the results on the same graph to see the final solution

OpenStudy (anonymous):

There is nothing to solve. I already have the graphs, and am working backwards to find the inequality.

OpenStudy (triciaal):

look at number 25

OpenStudy (triciaal):

x > -12 and x< -6

OpenStudy (freckles):

\[\text{ assume } a \text{ is positive } \\ |x-c| \le a \text{ means we are shading the interval }[ c-a , c+a] \\ |x-c| \ge a \text{ means we are shading everything outside the interval } (c-a,c+a)\]

OpenStudy (triciaal):